### v1

# Few-Shot Classification of Interactive Activities of Daily Living (InteractADL)

Abstract
Understanding Activities of Daily Living (ADLs) is a crucial step for different applications including assistive robots, smart homes, and healthcare. However, to date, few benchmarks and methods have focused on complex ADLs, especially those involving multi-person interactions in home environments. In this paper, we propose a new dataset and benchmark, InteractADL, for understanding complex ADLs that involve interaction between humans (and objects). Furthermore, complex ADLs occurring in home environments comprise a challenging long-tailed distribution due to the rarity of multi-person interactions, and pose fine-grained visual recognition tasks due to the presence of semantically and visually similar classes. To address these issues, we propose a novel method for fine-grained few-shot video classification called Name Tuning that enables greater semantic separability by learning optimal class name vectors. We show that Name Tuning can be combined with existing prompt tuning strategies to learn the entire input text (rather than only learning the prompt or class names) and demonstrate improved performance for few-shot classification on InteractADL and 4 other fine-grained visual classification benchmarks. For transparency and reproducibility, we release our code here.
1 Introduction
Figure 1: Our fine-grained visual recognition dataset InteractADL poses a challenging classification task for Visual Language Models (VLMs). Left: a sample timeline of multiple ADLs in our InteractADL dataset. We show two $3^{\text{rd}}$ person views along with annotated higher-level activity labels, temporal atomic actions, and dense spatiotemporal scene graphs. Right: dual encoder VLMs have a joint video-language embedding space, in which similarity scores between an input video and category names are computed. For these VLMs, we seek to use a small number of training examples to learn better class names that provide greater semantic separation and improved classification performance.
Activities of Daily Living (ADLs) are defined as a set of basic, essential, and routine tasks that most healthy people can perform without any assistance [Katz(1983)]. Recognizing ADLs is an essential component in several applications, from smart homes [Debes et al.(2016)Debes, Merentitis, Sukhanov, Niessen, Frangiadakis, and Bauer] and assistive robots [Lauretti et al.(2017)Lauretti, Cordella, Guglielmelli, and Zollo] to automated healthcare assessment [Desai et al.(2004)Desai, Grossberg, and Sheth] applications. Despite these essential use cases, there are few benchmarks and methods for understanding human ADLs at home, especially where there is interaction with other humans involved. Prior work has focused on ADLs at home with a single person present, such as Toyota Smart Home [Dai et al.(2022)Dai, Das, Sharma, Minciullo, Garattoni, Bremond, and Francesca] and Home Action Genome (HOMAGE) [Rai et al.(2021)Rai, Chen, Ji, Desai, Kozuka, Ishizaka, Adeli, and Niebles], or has presented benchmarks for ego-view scenes in specific rooms in the house, e.g., in the kitchen as in EPIC-Kitchens [Damen et al.(2018)Damen, Doughty, Farinella, Fidler, Furnari, Kazakos, Moltisanti, Munro, Perrett, Price, et al.].
There are several challenges associated with understanding ADLs at home. Specifically, (1) such activities are often long and sometimes take hours (e.g., house cleaning). They are often composed of several atomic actions (indivisible sequences of primitive actions that are completed without interruption and often take 2-3 seconds); (2) In addition to the relationships between humans and objects, it is crucial to model interactions between multiple humans when building models for recognizing ADLs. Specific interactions between humans are often rare, forming a long-tailed distribution that poses challenges for large-scale training; (3) ADLs consist of many classes that are visually and semantically similar, posing a fine-grained action recognition task that is challenging for many current methods.
In this paper, we introduce a new dataset and benchmark for complex video understanding of human interactions and ADLs in the home environment called InteractADL. The dataset includes RGB videos from multiple views, including each of two people’s ego-view and between 2-19 $3^{\text{rd}}$ person views per video. In addition to the object and human bounding boxes annotated in the dataset, it includes dense annotation of human-object (H-O) relationships and interactions between humans (H-H) in the form of spatiotemporal scene graphs. The sequences are further annotated with activity labels at two levels of abstraction, i.e., overall activity class and atomic action segments (see Fig. 1). The tasks for this benchmark comprise detecting dynamic H-H interactions as well as few-shot detection of higher-level activities and atomic actions in the videos.
Inspired by our new dataset and benchmark, we also propose a novel method for fine-tuning visual-language models (VLMs). Our method, called Name Tuning, enables greater semantic flexibility and separability between category names by allowing them to be fine-tuned for fine-grained recognition tasks. We show that Name Tuning achieves superior performance on InteractADL compared to existing dual encoder fine-tuning methods and show that Name Tuning achieves state-of-the-art performance across 3 separate few-shot video classification benchmarks and can perform on par with existing methods without using any video pre-training data.
2 Related Work
Activity Atomic Multi- Multi- View
Activity Datasets ADL Hours Classes Classes HOI HHI View person Type
ToyotaTrimmed [Das et al.(2019)Das, Dai, Koperski, Minciullo, Garattoni, Bremond, and Francesca] 53.1 5 31 M
TSU [Dai et al.(2022)Dai, Das, Sharma, Minciullo, Garattoni, Bremond, and Francesca] 188 5 51 M
CAD-120 0.5 10 20 S
DAHLIA [Vaquette et al.(2017)Vaquette, Orcesi, Lucat, and Achard] 33.4 7 - M
PKU-MMD [Liu et al.(2017)Liu, Hu, Li, Song, and Liu] 49 - 51 S
Charades [Sigurdsson et al.(2016)Sigurdsson, Varol, Wang, Farhadi, Laptev, and Gupta] 82 - 157 M
Charades-ego [Sigurdsson et al.(2018)Sigurdsson, Gupta, Schmid, Farhadi, and Alahari] 68.8 - 157 - E
ETRI-Activity3D [Jang et al.(2020)Jang, Kim, Park, Jang, Lee, and Kim] - - 55 S
MSR-DA3D [Wang et al.(2012)Wang, Liu, Wu, and Yuan] 16 - - S
RGBD-HuDaAct [Ni et al.(2011)Ni, Wang, and Moulin] 46 - 12 S
ActivityNet [Caba Heilbron et al.(2015)Caba Heilbron, Escorcia, Ghanem, and Carlos Niebles] 849 - 200 S
Kinetics-700 [Carreira et al.(2019)Carreira, Noland, Hillier, and Zisserman] 1167 - 700 S
AVA [Gu et al.(2018)Gu, Sun, Ross, Vondrick, Pantofaru, Li, Vijayanarasimhan, Toderici, Ricco, Sukthankar, et al.] 107.5 - 80 S
EPIC-Kitchens [Damen et al.(2018)Damen, Doughty, Farinella, Fidler, Furnari, Kazakos, Moltisanti, Munro, Perrett, Price, et al.] 55 - - E
MMAct [Kong et al.(2019)Kong, Wu, Deng, Klinkigt, Tong, and Murakami] 126.7 - 37 S
ActionGenome [Ji et al.(2020)Ji, Krishna, Fei-Fei, and Niebles] 82 - 157 M
Breakfast [Kuehne et al.(2014)Kuehne, Arslan, and Serre] 77 - 10 S
Ego4D [Grauman et al.(2021)Grauman, Westbury, Byrne, Chavis, Furnari, Girdhar, Hamburger, Jiang, Liu, Liu, et al.] 3670 - - E
HOMAGE [Rai et al.(2021)Rai, Chen, Ji, Desai, Kozuka, Ishizaka, Adeli, and Niebles] 25.4 - - E + S
InteractADL 242 50 334 M + T + E
Table 1: Comparison with other ADL (Activities of Daily Living) Datasets (top) and related video datasets (bottom). HHI: Human-human interaction, HOI: Human-object interaction. View types: M: Monitoring, S: Shooting, E: Ego, T: Top-Down.
Data Acquisition of ADLs at Home. Prior work in building datasets of human actions has varied in their degree of spontaneity, with some designs requiring participants to act out specific action classes and steps, such as PKU-MMD [Liu et al.(2017)Liu, Hu, Li, Song, and Liu], Charades [Sigurdsson et al.(2018)Sigurdsson, Gupta, Schmid, Farhadi, and Alahari], ETRI-activity3D [Jang et al.(2020)Jang, Kim, Park, Jang, Lee, and Kim], Rgbd-hudaact [Ni et al.(2011)Ni, Wang, and Moulin], MMAct [Kong et al.(2019)Kong, Wu, Deng, Klinkigt, Tong, and Murakami], and ActionGenome [Ji et al.(2020)Ji, Krishna, Fei-Fei, and Niebles]. In contrast, some other datasets have given more agency to actors with unscripted actions, including Toyota [Das et al.(2019)Das, Dai, Koperski, Minciullo, Garattoni, Bremond, and Francesca], DAHLIA [Vaquette et al.(2017)Vaquette, Orcesi, Lucat, and Achard], and Ego4D [Grauman et al.(2021)Grauman, Westbury, Byrne, Chavis, Furnari, Girdhar, Hamburger, Jiang, Liu, Liu, et al.].
Capturing multiple viewpoints can allow models to become more robust to variations of input. A few prior works have, accordingly, focused on acquiring multiple viewpoints, with [Das et al.(2019)Das, Dai, Koperski, Minciullo, Garattoni, Bremond, and Francesca] utilizing 7 cameras, [Vaquette et al.(2017)Vaquette, Orcesi, Lucat, and Achard] utilizing 3 Kinect v2 sensors, [Kong et al.(2019)Kong, Wu, Deng, Klinkigt, Tong, and Murakami] utilizing 4 camera views and 1 egocentric view, and [Kuehne et al.(2014)Kuehne, Arslan, and Serre] utilizing 3 to 5 views depending on the location.
Previous data acquisition methods from home environments often specialize in single-person actions, like Epic-Kitchens [Damen et al.(2018)Damen, Doughty, Farinella, Fidler, Furnari, Kazakos, Moltisanti, Munro, Perrett, Price, et al.]. Additionally, prior work has also focused on human-object interaction, namely JRDB-Act [Ehsanpour et al.(2021)Ehsanpour, Saleh, Savarese, Reid, and Rezatofighi] and ActionGenome [Ji et al.(2020)Ji, Krishna, Fei-Fei, and Niebles]. However, actions in the real world often involve multiple individuals interacting with each other and objects. Few datasets involve multiple individuals for a given task, and such configurations tend to comprise only a portion of the dataset [Liu et al.(2017)Liu, Hu, Li, Song, and Liu, Ji et al.(2020)Ji, Krishna, Fei-Fei, and Niebles]. In contrast to previous datasets, InteractADL captures spontaneous ADLs between two individuals and provides high-level action, atomic action, and spatiotemporal scene graph annotations to capture human-human interactions while providing both ego and $3^{\text{rd}}$ person views. We highlight our explicit differences with existing datasets in Table 1.
Large Pre-trained Visual-language Models. An increasingly popular paradigm when training visual-language systems is to leverage vast amounts of task-agnostic general internet data and use self-supervised learning to train large visual-language models (VLMs) that perform well on a variety of downstream tasks [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever, Alayrac et al.(2022)Alayrac, Donahue, Luc, Miech, Barr, Hasson, Lenc, Mensch, Millican, Reynolds, et al.]. Using matching visual-text pairs from the internet, such models can be trained with a variety of self-supervised objectives, including contrastive learning or image captioning [Yu et al.(2022)Yu, Wang, Vasudevan, Yeung, Seyedhosseini, and Wu]. By modeling computer vision tasks using natural language, these VLMs tend to exhibit powerful zero-shot and few-shot capabilities over a variety of domains [Zhou et al.(2021)Zhou, Yang, Loy, and Liu, Alayrac et al.(2022)Alayrac, Donahue, Luc, Miech, Barr, Hasson, Lenc, Mensch, Millican, Reynolds, et al.].
Fine-tuning Large Pre-trained Models. Fine-tuning pre-trained large language models (LLMs) for downstream tasks has been a very popular approach in Natural Language Processing (NLP) [He et al.(2021)He, Zhou, Ma, Berg-Kirkpatrick, and Neubig, Liu et al.(2021)Liu, Yuan, Fu, Jiang, Hayashi, and Neubig]. One approach for tuning large language models uses prompt-tuning methods to train token embeddings that act as a prompt alongside additional text input [Li and Liang(2021), Liu et al.(2021)Liu, Yuan, Fu, Jiang, Hayashi, and Neubig]. Versions of prompt-tuning [Zhou et al.(2021)Zhou, Yang, Loy, and Liu, Lu et al.(2022)Lu, Liu, Zhang, Liu, and Tian, Zhou et al.(2022)Zhou, Yang, Loy, and Liu] have been applied to CLIP, allowing efficient fine-tuning for image classification and text-image retrieval. We note that these existing methods use class names as fixed, unlearnable tokens within class-specific prompts, whereas our work demonstrates the potential advantages of fine-tuning these class name tokens. We highlight the difference between our method and existing prompt tuning methods in Fig. 6.
3 InteractADL
InteractADL is a multi-view RGB video dataset that provides detailed annotations for understanding daily living activities and interactions in a home environment. The dataset is recorded on sensors capturing video recordings synchronized across multiple viewpoints (ego-view and $3^{\text{rd}}$ person). Given the difficulty of obtaining videos of home environments, this dataset provides an ego-view and $3^{\text{rd}}$ person views for all videos. InteractADL provides hierarchical action labels with high-level activity and low-level action information, human and object bounding boxes, and dense human-human and human-object interaction labels.
3.1 Annotation Protocol
For each task performed in a video, pairs of participants were selected such that the same pairs were not used for different tasks. Vidat [Zhang et al.(2020)Zhang, Gould, and Ben-Shabat] was used as a tool to annotate activities and atomic actions per frame using object bounding boxes and semantic and instance regions. There were a total of 13 participants in the dataset recordings. The scene graph itself is annotated using a customized version of CVAT [Sekachev et al.(2020)Sekachev, Manovich, Zhiltsov, Zhavoronkov, Kalinin, Hoff, TOsmanov, Kruchinin, Zankevich, DmitriySidnev, Markelov, Johannes222, Chenuet, a andre, telenachos, Melnikov, Kim, Ilouz, Glazov, Priya4607, Tehrani, Jeong, Skubriev, Yonekura, vugia truong, zliang7, lizhming, and Truong], which similarly uses a bounding box approach for object detection and annotation.
In total, there are 26 cameras, including 2 ego-view cameras, and 24 $3^{\text{rd}}$ person view cameras to provide a variety of views. Within the $3^{\text{rd}}$ person views, some video angles comprise a viewpoint looking down from above (ceiling view), and we henceforth count these as part of the $3^{\text{rd}}$ person views. For each long-term video, both participants’ ego-view videos are provided, and a subset of the 3rd-person views are provided. Each long-term video contains at least two $3^{\text{rd}}$ person views, with a maximum of 19 $3^{\text{rd}}$ person videos.
3.2 Annotation Details
InteractADL contains (1) high-level activity labels, (2) atomic action class labels, (3) bounding boxes, and (4) spatiotemporal scene-graph labels. High-level activities encompass many atomic actions, sometimes denoted sub-activity labels in other datasets; for example, a high-level activity of unloading washing machine includes the atomic activities holding basket, holding detergent, etc. We provide temporally localized labels for 50 activity and 334 atomic actions. InteractADL contains in total 396 activity instances with 5,544 atomic action instances. Bounding boxes for both people and objects are also annotated, totaling 70 classes of bounding boxes (e.g., person, basket, clothes), with 105,697 annotated bounding boxes, of which 3,314 are humans and 102,383 are objects. Finally, we provide densely-annotated scene-graph labels including human-human and human-object interaction labels across 30 different relationship classes. In total, 224,449 relationships are annotated.
Annotated video sequences within InteractADL are quite long, averaging about 10.6 minutes. These long sequence durations allow for complex scene understanding, including action sequences comprised of several ordered atomic actions or longer activities. Atomic actions in InteractADL vary widely in length. While atomic actions are generally short (average of 9.77 seconds), some are as long as 654.9 seconds. We compare the annotations of InteractADL with existing datasets in Table 1. We note that in our analysis for this work, we focus on atomic action classification from a single view (ego, monitoring, or top-down).
4 Method
Figure 2: Comparison of input-text optimization methods for dual encoder VLMs. We contrast the methods presented in this work, Name Tuning and CoNa, with standard prompt engineering and prompt tuning (CoOp). We introduce learnable offset vectors to fine-tune class names in the input text.
Our approach seeks to allow for model-guided separation between semantically similar class names, which are common in fine-grained visual classification datasets. Similar class names, which are prevalent when actions happen in constrained environments like in InteractADL, can be challenging for many low-shot classification algorithms designed for VLMs. This is because they often rely on using the given class names from the dataset metadata [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever, Zhou et al.(2021)Zhou, Yang, Loy, and Liu, Zhang et al.(2021)Zhang, Fang, Zhang, Gao, Li, Dai, Qiao, and Li, Lu et al.(2022)Lu, Liu, Zhang, Liu, and Tian, Ju et al.(2022)Ju, Han, Zheng, Zhang, and Xie]. We hypothesize that the default class names for a dataset are likely not the most discriminative for a pre-trained VLM, since they were originally developed for human dataset curation and annotation rather than for model classification. In a similar way that prompt tuning can be used to improve upon hand-crafted prompts, we show that name tuning improves the given class names for a dataset to make them more discriminative.
4.1 Method Background and Notation
Zero-shot Video Classification. In our work, we focus on the sub-class of VLMs that possess a joint visual-language embedding space (dual encoders), and thus each VLM has a separate visual and text encoder, which we denote as $E^{v}$ and $E^{t}$, respectively. We denote $x$ as the input image or video, and $n_{i}\in\mathbb{R}^{l_{i}\times d_{\text{token}}}$ as the sequence of token embeddings produced by tokenizing the $i^{\text{th}}$ class’s name text. In the zero-shot setting, contrastive VLMs perform visual classification by computing the similarity of a visual embedding $E^{v}(x)$ with text embedding $E^{t}(n_{i})$ for each class name. Usually, a fixed prompt, denoted $q$, is prepended to the class name resulting in text embeddings $E^{t}([q,n_{i}])$. The VLM’s predicted class probabilities are therefore the softmax of visual-text embedding similarity for each class:
$p(y=i\mid x)=\frac{\exp(\frac{1}{\tau}\left<E^{t}([q,n_{i}]),E^{v}(x)\right>)}{\sum_{j=1}^{N}\exp(\frac{1}{\tau}\left<E^{t}([q,n_{j}]),E^{v}(x)\right>)},$ (1)
[…]
Prompt Tuning. Context-Optimization (abbreviated CoOp) was introduced as a method for extending CLIP from zero-shot image classification to few-shot image classification via prompt tuning [Zhou et al.(2021)Zhou, Yang, Loy, and Liu]. The method seeks to improve upon prompt engineering approaches by using learnable, fixed-size context vectors instead of using the natural language prompt $q$.
For CoOp, a context vector sequence $c^{\theta}\in\mathbb{R}^{l_{\text{context}}\times d_{\text{token}}}$ is prepended to every tokenized class name $n_{i}$, producing text embeddings $E^{t}([c^{\theta},n_{i}])$. Given a training dataset, the cross-entropy loss is backpropagated through the frozen text encoder to train the context vectors.
Dataset Prompt Tuning Name Tuning $\Delta$
MOMA Activities [Luo et al.(2022)Luo, Durante, Li, Xie, Liu, Jin, Huang, Li, Wu, Niebles, et al.] 95.7 97.9 +2.2
Kinetics-100 [Kay et al.(2017)Kay, Carreira, Simonyan, Zhang, Hillier, Vijayanarasimhan, Viola, Green, Back, Natsev, Suleyman, and Zisserman] 92.7 94.7 +2.0
MOMA Sub-activities [Luo et al.(2022)Luo, Durante, Li, Xie, Liu, Jin, Huang, Li, Wu, Niebles, et al.] 76.6 78.2 +1.6
Table 2: Comparison of text input tuning methods across few-shot activity recognition benchmarks. We show show that our proposed method, name tuning, consistently outperforms prompt tuning for 5-shot 5-way classification on Kinetics and MOMA-LRG using a CLIP ViT-B32 backbone. Results are averaged across 20 runs with random seeds.
4.2 Name Tuning
Our method, Name Tuning, seeks to improve the expressiveness of provided class names directly, by training them to better match the content of example videos. For each class, we train a class name offset $\epsilon_{i}^{\theta}\in\mathbb{R}^{l_{\text{class}}\times d_{\text{token}}}$, matching the shape of the class name’s token embedding sequence $n_{i}$ and initialized to zero. Each class name offset is added to the corresponding class name, producing text embeddings $E^{t}([q,n_{i}+\epsilon_{i}^{\theta}])$.
Like CoOp, we train these parameters using a few-shot training dataset, backpropagating through the frozen text encoder to update the name offset parameters. To avoid overfitting when given only a few training examples per class, we also include an $L_{2}$ regularization penalty for the name offsets, which constrains the tuned class names to stay near the original class names, retaining the prior information they convey, resulting in the following loss function:
$J(\theta,\mathcal{D})=-\sum_{x,y\sim D}y\cdot\log p(y\mid x)+\alpha\frac{1}{2}\sum_{i=1}^{N}||\epsilon_{i}^{\theta}||_{2}^{2},$ (2)
where $||\cdot||_{2}$ denotes the $L_{2}$ norm, and $\alpha$ is a hyperparameter that corresponds to the strength of regularization. We search over $\alpha\in\{0.01,0.1,1.0,10.0\}$ (see Appendix for more details).
4.3 CoNa
Name Tuning is entirely compatible with the existing CoOp method since it tunes separate parts of the input text, and we observe that the combination of the two (abbreviated CoNa) can boost performance on some datasets. CoNa jointly trains a class-shared context $c^{\theta}\in\mathbb{R}^{l_{\text{context}}\times d_{\text{token}}}$ and class-specific name offsets $\epsilon_{i}^{\theta}\in\mathbb{R}^{l_{\text{class}}\times d_{\text{token}}}$, producing text embeddings $E^{t}([c^{\theta},n_{i}+\epsilon_{i}^{\theta}])$.
[…]
4.4 Few-shot Learning Paradigm
[…]
4.5 Datasets and Data Splits
[…]
5 Experiments and Results
[…]
Method 1 shot 2 shots 4 shots 8 shots
Linear Probe 3.95 4.94 5.41 6.71
VL-Prototype 3.83 5.02 4.94 5.81
CoOp 3.40 4.98 5.92 5.85
VIFI-CLIP* 3.36 3.88 5.50 5.97
Name Tuning 3.87 4.74 5.29 5.77
CoNa 3.87 5.21 6.00 6.95
Table 3: 56-way few-shot classification results on the InteractADL dataset. We compare our proposed methods (Name Tuning and CoNa) with existing methods for finetuning VLMs for visual classification. For fair comparison, we only include text-branch tuning for ViFi-CLIP. Best results in each column are typeset in bold.
5.1 Results on Other Activity Recognition Datasets
In addition to showing results for InteractADL, we demonstrate that Name Tuning and CoNa, without any video pretraining, set a new state-of-the-art for Kinetics (Table 4), and MOMA-LRG (Table 5) and provide strong few-shot performance on UCF-101 (Table 6). For the meta-learning splits, we compare against the state-of-the-art meta-learning (vision-only) baselines on the standard 5-shot, 5-way classification task (evaluating unseen classes from the meta-test set). Our VLM-based method does not make use of the meta-train set, and we choose hyperparameters based on 5-shot 5-way performance on the meta-validation set. Since we do not have a held-out validation set, we do not perform model checkpointing, and simply use the final-epoch model for validation and test results. We report the average test-set accuracy from 20 random seeds. In general, we find that our VLM-based method works well on these meta-learning activity recognition datasets and outperforms the vision-only few-shot approaches by a significant margin without requiring a meta-training set.
Method Meta-Training Free Accuracy
OTAM [Cao et al.(2020)Cao, Ji, Cao, Chang, and Niebles] 85.8
AMeFu-Net [Fu et al.(2020)Fu, Zhang, Wang, Fu, and Jiang] 86.8
STRM [Thatipelli et al.(2022)Thatipelli, Narayan, Khan, Anwer, Khan, and Ghanem] 86.7
MTFAN [Wu et al.(2022)Wu, Zhang, Zhang, Wu, and Zhang] 87.4
Name Tuning (ours) 94.7
CoNa (ours) 92.0
Table 4: 5-shot 5-way few-shot classification results for Kinetics. We compare our proposed methods (Name Tuning and CoNa) with results from previous approaches that rely upon a large meta-training set. Despite not using the meta-training set, we achieve state-of-the-art results for few-shot activity recognition on Kinetics. We use the meta-validation set for hyperparameter tuning.
Method MT Free Activity Acc. Subactivity Acc.
CMN [Zhu and Yang(2018)] 86.3 66.6
OTAM [Cao et al.(2020)Cao, Ji, Cao, Chang, and Niebles] 92.1 72.6
VideoCLIP-GVLM [Luo et al.(2022)Luo, Durante, Li, Xie, Liu, Jin, Huang, Li, Wu, Niebles, et al.] 84.8 32.7
Frozen-GVLM [Luo et al.(2022)Luo, Durante, Li, Xie, Liu, Jin, Huang, Li, Wu, Niebles, et al.] 92.5 26.3
Name Tuning (ours) 97.9 78.2
CoNa (ours) 95.7 76.4
Table 5: 5-shot 5-way few-shot classification results for MOMA-LRG. We compare our proposed methods (Name Tuning and CoNa) with results from previous meta-learning works. By leveraging pre-trained VLMs, we significantly outperform existing methods and do not require a large meta-training set. We use the meta-validation set for hyperparameter tuning (MT Free = Meta-Training Free).
Method 1 shot 2 shots 4 shots 8 shots
Linear Probe 76.8 84.5 90.1 92.5
VL-Prototype 79.4 86.2 91.0 92.8
CoOp 85.3 87.1 89.8 90.8
ViFi-CLIP* 85.3 88.1 91.5 92.6
Name Tuning 86.3 87.5 89.1 92.7
CoNa 85.5 88.0 91.2 93.5
Table 6: 101-way few-shot classification results for UCF-101. We compare our proposed methods (Name Tuning and CoNa) with existing methods for finetuning VLMs for visual classification. For fair comparison, we only include text-branch tuning for ViFi-CLIP. We note that ViFi-CLIP requires extensive video pretraining on Kinetics-400 and achieves similar results to CoNa that only uses averaged frame-level CLIP embeddings. We bold the highest accuracy for each number of shots.
6 Discussion
Comparison of Text Tuning Strategies.
In Table 2, we show the average difference in performance between traditional prompt tuning and name tuning across three few-shot activity recognition benchmarks. We note that the main requisite of CoOp is the existence of reliable and distinguishable class names, whereas Name Tuning requires a good baseline prompt to use to learn class names (for our results, we use “a video of {}").
Comparison with Meta-learning Methods.
In our experiments, we find that our VLM-based few-shot learning method outperforms the current state-of-the-art meta-learning approaches on all benchmarked datasets. This is notable since the VLM-based methods do not require a large meta-training set derived from visually similar classes and can instead leverage broad pre-training from internet scale data.
Performance on InteractADL.
In Table 3, we show that InteractADL is a very challenging dataset where the benchmarked VLM-based methods perform significantly worse (3-7%) than other activity recognition datasets (e.g., 76-94% for UCF-101). We hypothesize there are several reasons for this gap, including the domain gap between VLM pre-training data and home environment settings (and ego-view data), along with the fine-grained nature and subtle visual distinctions between atomic actions in our dataset. Regardless, due to the challenging nature of few-shot atomic action classification in home environments (with similar backgrounds) and the rich semantic annotations of InteractADL, we believe that InteractADL provides a challenging benchmark for fine-grained activity recognition at home.
7 Conclusion
In this paper, we introduced a new dataset, InteractADL, which includes multi-view data of multiple humans performing ADLs at home. To tackle the few-shot activity recognition problem posed by this dataset, we introduce a novel method, Name Tuning, for learning model-specific class names that expands upon existing prompt tuning strategies. Through this method, we also achieve a new state-of-the-art in three different few-shot activity recognition benchmarks (MOMA-Activities, MOMA-Subactivities, and Kinetics). We demonstrate the challenges for adapting pre-trained VLMs for our fine-grained video dataset, despite their impressive zero-shot performance on relatively high-level datasets. For transparency, reproducibility, and to accelerate research in this area, we release our code at the following URL.
Appendix
Figure 3: We detail the floorplan of the home environment used to collect data for InteractADL. InteractADL is recorded in a real home environment with sensors (represented in red) placed throughout the many rooms to provide a subset of views for each long-term video, including ego-view, $3^{\text{rd}}$ person, and top-down. Ego-view is provided for all videos, and at least two $3^{\text{rd}}$ person views (including ceiling views) are provided for each video.
Appendix A InteractADL Details
Floorplan We show a detailed floorplan of a real home that was used for collecting data for InteractADL in Figure 3. Dataset Samples We show three example frames from our labeled dataset in Figure 4.
Figure 4: We show three example scene graph annotations from InteractADL corresponding to three separate rooms in the home. Our annotations cover the primary household objects and actors in the scene and interactions relevant for ADLs.
Appendix B Name Tuning Ablation and Analysis
A majority of our ablations used either the Kinetics or MOMA-LRG datasets. We chose these datasets since (1) Kinetics is a widely used benchmark for video understanding and activity recognition, and (2) MOMA-LRG was a dataset specifically constructed for challenging VLMs to require more fine-grained understanding of video scenes.
Effect of Prompt Choice. To help better understand the effect of prompt choice on the performance of Name Tuning, we show the performance of three different prompts on MOMA Sub-activities in Figure 5. We also show how varying the prompt for Name Tuning can effect 5-shot all-way performance on UCF-101 and Kinetics in Figure 7. For this analysis, to better observe the effects of prompt choice and disentangle other hyperparameter choices, we fix the learning rate to $10^{-4}$ and the name regularization parameter to 1 and thus do not perform any hyperparameter search, resulting in slightly worse performance than what we report in the main paper. For Figure 7, we use the 28 prompts from [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever] that were originally used to obtain zero-shot performance for the UCF-101 dataset. We show results for each prompt averaged across 20 runs. To summarize our findings: good prompt choice can result in small improvements in performance (up to 2% absolute). For the exact prompts used, see the prompt templates described in https://github.com/openai/CLIP/blob/main/data/prompts.md#ucf101.
Figure 5: Name Tuning performance on the MOMA Sub-activities dataset using various prompts and CLIP as a pre-trained backbone. We label the prompt besides Name Tuning (NT) in the legend and show linear probe and CoOp as baselines.
Randomly Initialized Class Names. We also explored how important the initialization of class names is for Name Tuning performance. Thus, we test the performance on MOMA Sub-activities with class names randomly initialized to $\mathcal{N}(0,0.02)$, following the initialization procedure in CoOp [Zhou et al.(2021)Zhou, Yang, Loy, and Liu]. Since we do not use a semantically meaningful initialization for class names, we do not use the regularization term to penalize large learned offset vectors, thus allowing greater exploration of the input vector space. We show that random initialization under-performs initializing input vectors to the default dataset class names and report results for 1, 2, 4, 8, and 16 shots in Figure 6.
Figure 6: Performance on the MOMA Sub-activities dataset for various VLM fine-tuning methods using CLIP as a pre-trained backbone. For Name Tuning (NT), we use "A video of {}". We show that using randomly initialized (rand init) class names under-performs using the default dataset class names (class init) for Name Tuning.
[…]
Results on Image Classification Datasets We verify that Name Tuning can also perform well on fine-grained image classification datasets when using CLIP as our backbone. We select a subset of the datasets from [Zhou et al.(2021)Zhou, Yang, Loy, and Liu] that are and challenging for zero-shot CLIP and visually fine-grained. Namely, we use Oxford Flowers [Nilsback and Zisserman(2006)], Oxford Pets [Parkhi et al.(2012)Parkhi, Vedaldi, Zisserman, and Jawahar], and EuroSAT [Helber et al.(2019)Helber, Bischke, Dengel, and Borth]. Generally, Name Tuning performs similarly to CoOp for the datasets evaluated. We show results in Table 7.
Dataset Method 1-shot 2-shot 4-shot 8-shot 16-shot
Oxford Flowers CoOp [Zhou et al.(2021)Zhou, Yang, Loy, and Liu] 74.9 80.4 86.8 91.4 93.6
Name Tuning 76.2 83.4 89.0 92.4 95.0
Oxford Pets CoOp [Zhou et al.(2021)Zhou, Yang, Loy, and Liu] 87.1 86.4 87.8 89.6 90.1
Name Tuning 82.7 84.8 86.5 87.5 89.0
EuroSAT CoOp [Zhou et al.(2021)Zhou, Yang, Loy, and Liu] 54.4 59.8 73.8 80.6 83.9
Name Tuning 54.3 66.6 72.6 82.2 85.1
Table 7: Name Tuning and CoOp results on few-shot image classification. Both methods are evaluated using CLIP ViT-B32 as the pretrained VLM. Generally, Name Tuning performs similar to CoOp for fine-grained, few-shot image classification datasets with semantically similar classes.
B.1 Name Tuning Comparison to CoOp-CSC
CoOp-CSC is a version of the CoOp algorithm that trains class-specific context (CSC) embedding for each visual category. The authors find that it generally performs worse than the shared-context version of CoOp, though they note that it has potential in fine-grained environments [Zhou et al.(2021)Zhou, Yang, Loy, and Liu]. In Table 8, we compare CoOp-CSC against both CoOp and our algorithm CoNa for the fine-grained dataset MOMA Sub-activities. Our results show that shared-context with name-tuning (CoNa, ours) outperforms CoOp both with shared-context and class-specific-context training across most task settings in both datasets. Though the performance difference is large for low-shot settings, the algorithms tend to have closer performance in the 8 and 16-shot settings. Although both CoNa and CoOp-CSC have the ability to learn class-specific variables, CoOp-CSC must learn them from scratch whereas CoNa is given a helpful starting point by leveraging the dataset class names with an offset initialized to zero. CoOp-CSC also does not share information between classes, and thus can only learn class-specific variables, whereas CoNa uses shared context vectors. We hypothesize these factors may contribute to why CoNa performs better than CoOp-CSC in the few-shot setting.
Method 1-shot 2-shot 4-shot 8-shot 16-shot
CoOp 39.4, 37.2, 23.1 42.9, 30.3, 26.1 43.3, 38.3, 31.6 45.6, 40.4, 38.6 46.4, 43.6, 37.7
CoOp-CSC 31.5, 33.3, 26.5 34.9, 38.0, 31.1 37.4, 38.7, 33.8 40.5, 42.8, 38.7 44.7, 45.7, 40.9
Name Tuning 39.6, 36.5, 29.6 41.4, 38.8, 33.6 44.6, 40.5, 37.6 46.8, 43.3, 38.0 49.2, 46.3, 41.7
CoNa 41.9, 37.1, 29.2 44.2, 38.5, 32.6 46.5, 40.9, 36.9 47.5, 42.2, 38.7 49.7, 45.9, 40.9
Table 8: Comparison vs CoOp (CSC) on the MOMA-Subactivities activity recognition dataset. We show results using CLIP, MILES, and VideoCLIP by denoting each entry with three values representing the performance of the three VLMs, respectively. We bold the highest accuracy for each VLM, and for each number of shots.
Appendix C Results Across VLMs
Method 0-shot 1-shot 2-shot 4-shot 8-shot 16-shot
Prompt Engineering 81.3, 70.7, 49.5 81.3, 70.7, 49.5 81.3, 70.7, 49.5 81.3, 70.7, 49.5 81.3, 70.7, 49.5 81.3, 70.7, 49.5
VL-Prototype 85.8, 75.7, 60.7 87.5, 80.3, 63.4 85.5, 84.2, 65.4 88.4, 86.4, 67.9 90.5, 88.5, 69.7
Linear Probe 67.2, 70.6, 54.6 80.6, 77.9, 61.7 86.0, 81.7, 70.5 90.8, 87.8, 77.1 93.0, 89.6, 79.2
CoOp 88.7, 73.7, 59.0 90.9, 78.8, 61.4 91.7, 82.1, 66.8 93.6, 84.3, 73.1 95.5, 86.6, 77.7
Name Tuning (ours) 88.9, 78.4, 61.9 91.7, 79.8, 67.0 91.0, 84.6, 73.3 93.7, 85.9, 74.9 95.8, 87.8, 77.7
CoNa (ours) 87.9, 75.1, 58.8 91.3, 78.9, 63.9 92.3, 82.1, 70.7 94.0, 82.5, 71.8 94.7, 86.5, 78.8
Table 9: Low-shot classification results on the MOMA-LRG Activities dataset. We show results using CLIP, MILES, and VideoCLIP by denoting each entry with three values representing the performance of the three VLMs, respectively. We bold the highest accuracy for each VLM and for each number of shots.
Method 0-shot 1-shot 2-shot 4-shot 8-shot 16-shot
Prompt Engineering 37.9, 32.0, 23.8 37.9, 32.0, 23.8 37.9, 32.0, 23.8 37.9, 32.0, 23.8 37.9, 32.0, 23.8 37.9, 32.0, 23.8
VL-Prototype —- 40.0, 37.8, 27.9 40.8, 40.0, 30.6 41.5, 41.6, 32.7 41.6, 42.7, 34.3 42.3, 44.3, 36.7
Linear Probe 28.6, 31.0, 26.2 33.1, 35.2, 31.0 37.1, 38.9, 36.1 42.1, 42.0, 38.6 46.1, 44.8, 42.5
CoOp 39.4, 37.2, 23.1 42.9, 30.3, 26.1 43.3, 38.3, 31.6 45.6, 40.4, 38.6 46.4, 43.6, 37.7
Name Tuning (ours) 39.6, 36.5, 29.6 41.4, 38.8, 33.6 44.6, 40.5, 37.6 46.8, 43.3, 38.0 49.2, 46.3, 41.7
CoNa (ours) 41.9, 37.1, 29.2 44.2, 38.5, 32.6 46.5, 40.9, 36.9 47.5, 42.2, 38.7 49.7, 45.9, 40.9
Table 10: Low-shot classification results on the MOMA-LRG Sub-activities dataset. We show results using CLIP, MILES, and VideoCLIP by denoting each entry with three values representing the performance of the three VLMs, respectively. We bold the highest accuracy for each VLM and for each number of shots.
Method 0-shot 1-shot 2-shot 4-shot 8-shot 16-shot
Prompt Engineering 72.4, 54.6, 37.6 72.4, 54.6, 37.6 72.4, 54.6, 37.6 72.4, 54.6, 37.6 72.4, 54.6, 37.6 72.4, 54.6, 37.6
VL-Prototype 72.2, 59.3, 41.2 72.3, 61.9, 44.5 72.2, 64.2, 46.5 71.9, 66.6, 48.5 70.9, 68.2, 50.5
Linear Probe 37.0, 45.9, 28.7 49.8, 55.2, 38.2 58.3, 61.9, 48.3 65.9, 66.8, 54.9 71.1, 70.6, 57.8
CoOp 70.0, 59.1, 40.1 72.1, 60.7, 41.4 73.0, 63.4, 47.2 75.3, 66.6, 51.5 76.1, 68.2, 56.0
Name Tuning (ours) 73.2, 58.3, 42.3 73.8, 60.7, 45.6 74.7, 63.7, 49.0 76.0, 66.5, 53.3 77.7, 69.2, 57.4
CoNa (ours) 69.9, 59.0, 41.3 71.7, 61.5, 44.2 73.4, 63.7, 47.0 76.0, 66.8, 52.1 77.8, 68.7, 55.8
Table 11: Low-shot classification results on the Kinetics-100 dataset. We show results using CLIP, MILES, and VideoCLIP by denoting each entry with three values representing the performance of the three VLMs, respectively. We bold the highest accuracy for each VLM and for each number of shots.
We show results for 3 dual encoders VLMs in our experiments. Specifically, we try CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever], VideoCLIP [Xu et al.(2021)Xu, Ghosh, Huang, Okhonko, Aghajanyan, Metze, Zettlemoyer, and Feichtenhofer], and MILES [Ge et al.(2022)Ge, Ge, Liu, Wang, Wu, Shan, Qie, and Luo]. Results can be seen in Tables 9, 10, and 11, where our methods (Name Tuning and CoNa) generalize to VLMs beyond CLIP and perform well on all benchmarked datasets. CLIP outperforms MILES and VideoCLIP as a pre-trained VLM across all methods and datasets tested, which we found to be particularly notable since CLIP can only be used as a frame-level encoder. We also show results on the 5-shot, 5-way task for each VLM backbone in Tables 12, 13, and 14. Similarly, CLIP outperforms the other VLMs and Name Tuning outperforms all other benchmarked methods.
Method Meta-training Free? Accuracy
OTAM [Cao et al.(2020)Cao, Ji, Cao, Chang, and Niebles] 85.8
AMeFu-Net [Fu et al.(2020)Fu, Zhang, Wang, Fu, and Jiang] 86.8
VL-Prototype 92.9, 93.5, 85.0
Linear Probe 89.2, 92.7, 86.8
CoOp [Zhou et al.(2021)Zhou, Yang, Loy, and Liu] 92.7, 91.7, 86.7
Name Tuning (ours) 94.7, 90.8, 85.8
CoNa (ours) 92.0, 92.0, 86.5
Table 12: Comparison vs state-of-the-art meta-learning algorithms for the Kinetics dataset. We show performance using standard 5-shot 5-way accuracy on the meta-test set. We note that the VLM-based few-shot learning approaches do not require a meta-training set. For the VLM-based approaches, we show results using CLIP, MILES, and VideoCLIP by denoting each entry with three values representing the performance of the three VLMs, respectively. We bold the single best accuracy in the table.
Method Meta-training Free? Accuracy
CMN [Zhu and Yang(2018)] 86.3
OTAM [Cao et al.(2020)Cao, Ji, Cao, Chang, and Niebles] 92.07
VL-Prototype 86.5, 95.9, 92.3
Linear Probe 87.4, 96.1, 75.9
CoOp [Zhou et al.(2021)Zhou, Yang, Loy, and Liu] 88.6, 95.7, 92.8
Name Tuning (ours) 97.9, 95.6, 92.7
CoNa (ours) 90.0, 95.7, 92.9
Table 13: Comparison vs state-of-the-art meta-learning algorithms for the MOMA Activities dataset. We show performance using standard 5-shot 5-way accuracy on the meta-test set. We note that the VLM-based few-shot learning approaches do not require a meta-training set. For the VLM-based approaches, we show results using CLIP, MILES, and VideoCLIP by denoting each entry with three values representing the performance of the three VLMs, respectively. We bold the single best accuracy in the table.
Method Meta-training Free? Accuracy
CMN [Zhu and Yang(2018)] 66.6
OTAM [Cao et al.(2020)Cao, Ji, Cao, Chang, and Niebles] 72.6
VL-Prototype 71.2, 75.7, 75.6
Linear Probe 71.3, 75.1, 74.5
CoOp [Zhou et al.(2021)Zhou, Yang, Loy, and Liu] 74.0, 72.8, 76.6
Name Tuning (ours) 71.0, 74.0, 78.2
CoNa (ours) 71.3, 73.6, 76.4
Table 14: Comparison vs state-of-the-art meta-learning algorithms for the MOMA Sub-activities dataset. We show performance using standard 5-shot 5-way accuracy on the meta-test set. We note that the VLM-based few-shot learning approaches do not require a meta-training set. For the VLM-based approaches, we show results using CLIP, MILES, and VideoCLIP by denoting each entry with three values representing the performance of the three VLMs, respectively. We bold the single best accuracy in the table.
Appendix D Hyperparameter selection
To evaluate across all methods consistently, we choose the subset of hyperparameters outlined below that maximize accuracy on the validation set, and use that same set of hyperparameters for a new run on the training set, reporting resulting accuracy on the test set. We perform this individually for each combination of dataset, VLM backbone, method, and number of shots.
VL-Prototype. For the text weight, we sample 16 values from the interval $[10^{-2},10^{2}]$ log-uniformly. We also try two prompt ensembles, one is the set of prompts originally used by TIP-Adapter [Zhang et al.(2021)Zhang, Fang, Zhang, Gao, Li, Dai, Qiao, and Li], and the other is the highest performing set of prompts obtained from our own prompt-engineering. We report results averaged across 32 runs for the traditional data splits, and report results averaged across 1,000 runs for the meta-learning splits.
Linear Probe. We follow the approach used to run the linear probe baseline in the original CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever] paper and train a logistic regression classifier using L-BFGS with a maximum of 1,000 iterations. We determine $\lambda$, the L2 regularization strength coefficient by doing a hyperparameter sweep on the validation set by sampling 96 values from the interval $[10^{-6},10^{6}]$ log-uniformly. We report results averaged across 5 runs when using the traditional data splits, and report results averaged across 100 runs for the meta-learning splits.
CoOp For CoOp, we search over a context length of 8 and 16, and train for 50 epochs using SGD with batch size of 8. We perform a grid search over learning rates of $\{6.25\times 10^{-5},5\times 10^{-4},2\times 10^{-3},4\times 10^{-3}\}$. We checkpoint after each epoch and use the checkpoint with the best performance on the validation set.
Name Tuning. For Name Tuning, we use "A video of {}" as our context prompt. We train for 20 epochs using AdamW with batch size of 8. We perform a grid search over learning rates of $\{1\times 10^{-5},1\times 10^{-4},4\times 10^{-3},1\times 10^{-3}\}$, and $\{0.01,0.1,1,10\}$ for the name regularization hyperparameter, $\alpha$. We checkpoint after each epoch and use the checkpoint with the best performance on the validation set.
CoNa. For CoNa, we use a fixed learning rate of $10^{-5}$, and train for 20 epochs using AdamW with batch size of 8. We perform grid search over the regularization coefficient $\alpha\in\{1,5,10,20\}$ and use a context length of 4. We checkpoint after each epoch and use the checkpoint with the best performance on the validation set.